OpenStudy (anonymous):

Water is dripping out from a conical funnel at a uniform rate of 2 cm3/s through a tiny hole at the vertex in the bottom. When the slant height of the water is 4 cm, find the rate of decrease of the slant height of the water-cone, given that the vertical angle of the funnel is 120

OpenStudy (anonymous):

@hartnn @experimentX

OpenStudy (bahrom7893):

Figure out the formula for the slant height of a cone. Then differentiate it with respect to t, plug in all the given information and you'll have your answer.

OpenStudy (bahrom7893):

Maybe experimentX can elaborate, but I'm really busy at the moment.

OpenStudy (anonymous):

Radius remains Constant ryt?

OpenStudy (bahrom7893):

No radius changes as well since it's a cone.

OpenStudy (anonymous):

V = pie*r^2*h/3 2r = h r = h/2 V = pie*h^3/12 dV/dt = pie /12 * 3 h^2 dh/dt

OpenStudy (anonymous):

dl/dt = 5/pie...am i correct?

OpenStudy (bahrom7893):